# A multiplicity formula for the Milnor number of smoothable curves

### v2

@@ Abstract.
We derive a multiplicity formula for the Milnor number of a reduced smoothable curve singularity generalizing a well-known formula due to Lê, Greuel and Teissier for complete intersection curves. We obtain a multiplicity characterization of Whitney equisingularity for families of locally smoothable curves.
@@
Milnor number, Euler characteristic, delta invariant, complete intersection discrepancy, conormal space, polar varieties, Hilbert–Samuel and Buchsbaum–Rim multiplicities, Whitney equisingularity
@@ 1. Introduction
Let $(X_{0},0)\subset(\mathbb{C}^{n},0)$ be a reduced complex analytic curve singularity. In [BG80] Buchweitz and Greuel define the Milnor number of $(X_{0},0)$ as follows. Let $n\colon\overline{(X_{0},0)}\rightarrow(X_{0},0)$ be the normalization morphism. Denote by $\omega_{X_{0},0}$ the dualizing module of Grothendieck. Denote by $d$ the composition of $\Omega_{X_{0},0}^{1}\rightarrow n_{*}\Omega_{\overline{X_{0},0}}^{1}\cong n_{*}\omega_{\overline{X_{0},0}}\rightarrow\omega_{X_{0},0}$ and the exterior derivation $\mathcal{O}_{X_{0},0}\rightarrow\Omega_{X_{0},0}^{1}$. Then the Milnor number $\mu$ of $(X_{0},0)$ is defined as
@@
In [BG80, Prp. 1.2.1] Buchweitz and Greuel showed that $\mu=2\delta+r-1$, where $\delta:=\dim_{\mathbb{C}}(\overline{\mathcal{O}_{X_{0},0}}/\mathcal{O}_{X_{0},0})$ and $r$ is the number of branches of $(X_{0},0)$. This is a formula discovered for plane curves by Milnor. When $(X_{0},0)$ is smoothable, Bassein [Bas77] proved that $\mu$ equals the first Betti number of the smoothing; in particular, for complete intersection curves $\mu$ coincides with the usual Milnor number.
@@
We say that $(X_{0},0)$ is smoothable if there exists a flat deformation $h\colon(X,0)\rightarrow(Y,0)$ with $Y$ smooth of dimension one such that $(h^{-1}(0),0)=(X_{0},0)$ and the fiber $h^{-1}(y):=X_{y}$ is smooth for $y\neq 0$ close enough to 00. The main result of this paper generalizes the formula above to the case of smoothable curves by introducing a correction term that loosely speaking measures how far $(X_{0},0)$ is from a complete intersection. Let $(Z_{0},0)\subset(\mathbb{C}^{n},0)$ be a generic complete intersection curve which contains $(X_{0},0)$ as a subvariety. Set $W_{0}:=\overline{Z_{0}\setminus X_{0}}$. Denote by $I_{0}(X_{0},W_{0})$ the intersection number of $X_{0}$ and $W_{0}$ at 00 (see [Hir57, pg. 179]). It is equal to $\dim_{\mathbb{C}}\mathcal{O}_{\mathbb{C}^{n},0}/(I_{X_{0}}+I_{W_{0}})$, where $I_{X_{0}}$ and $I_{W_{0}}$ are the ideals of $X_{0}$ and $W_{0}$ in $\mathcal{O}_{\mathbb{C}^{n},0}$, respectively. We will call $I_{0}(X_{0},W_{0})$ the complete intersection discrepancy of $X_{0}$.
@@
When $(X_{0},0)$ is a complete intersection, then $I_{W_{0}}=(1)$ and so $I_{0}(X_{0},W_{0})$ vanishes. Thus (1) recovers the Lê–Greuel–Teissier formula.
@@
Whitney equisingularity. Throughout the paper when dealing with complex analytic germs we will work with suitably small representatives of these germs. Let $(X,0)\subset(\mathbb{C}^{n+k},0)$ be a reduced complex analytic space of pure dimension $k+1$, and let $(Y,0)$ be a linear subspace of $(X,0)$ of dimension $k$. Choose an embedding of $(X,0)$ in $\mathbb{C}^{n+k}=\mathbb{C}^{n}\times\mathbb{C}^{k}$, so that $(Y,0)$ is represented by $0\times U$, where $U$ is an open neighborhood of 00 in $\mathbb{C}^{k}$. Let $\mathrm{pr}:\mathbb{C}^{n}\times\mathbb{C}^{k}\rightarrow\mathbb{C}^{k}$ be the projection on the second factor and let $i:\mathbb{C}^{k}\rightarrow\mathbb{C}^{n}\times\mathbb{C}^{k}$ be the inclusion $i(y)=(0,y)$. Denote by $\pi_{Y}$ be the retraction $i\circ\mathrm{pr}_{|X}$. View $X$ as the total space of the family $\pi_{Y}:(X,0)\rightarrow(Y,0)$. For each $y\in Y$ set $X_{y}:=X\cap\pi^{-1}_{Y}(y)$. We say $H$ is a tangent hyperplane at $x\in X-Y$ if $H$ is a hyperplane in $\mathbb{C}^{n+k}$ that contains the tangent space $T_{x}X$. Let $(x_{i})$ be a sequence of points from $X-Y$ and $(y_{i})$ be a sequence of points from $Y$ both converging to 00. Suppose that the sequence of secants $(\overline{x_{i}y_{i}})$ has limit $l$ and the sequence of tangent hyperplanes $\{T_{x_{i}}X\}$ has limit $T$. We say that $(X,0)\rightarrow(Y,0)$ is Whitney equisingular at 00, or that the pair of strata $(X-Y,Y)$ satisfies the Whitney conditions at 00 if $l\subset T$.
@@ 2.1. Conormal spaces
Denote $\mathrm{Image}(\delta^{*})$ by $J_{\mathrm{rel}}(X)$ and call it the relative Jacobian module of $X$. It’s a module contained in the free module $\mathcal{O}_{X,0}^{p}$. Define the Rees algebra $\mathcal{R}(J_{\mathrm{rel}}(X))$ to be the subalgebra of $\mathrm{Sym}(\mathcal{O}_{X,0}^{p})$ generated by $J_{\mathrm{rel}}(X)$ placed in degree $1$. Define the relative conormal space $C_{\mathrm{rel}}(X)$ of $X$ to be the closure in $X\times\check{\mathbb{P}}^{n-1}$ of the set of pairs $(x,H)$ where $x$ is a smooth point of $X_{h^{-1}(h(x))}$ and $H$ is a (tangent) hyperplane in $\mathbb{C}^{n}$ at $x$ containing $T_{x}X_{h^{-1}(h(x))}$. We have
@@ 2.3. Multiplicities
Let $(V,0)$ be a reduced complex analytic variety. Let $M$ be an $\mathcal{O}_{V}$-module such that its rank at the generic point of each irreducible component of $V$ is $e$. Assume $M$ is contained in a free module $\mathcal{O}_{V}^{p}$ for some $p$. Denote by $[M]\in\mathrm{Mat}(p\times c,\mathcal{O}_{V})$ a presentation matrix of $\mathcal{O}_{V}^{p}/M$. Let $A\in\text{Mat}(e\times p,\mathbb{C})$ be a generic matrix such that $[M_{e}]:=A[M]$ has rank $e$ at the generic point of each irreducible component of $(V,0)$. Let $M_{e}$ be the $\mathcal{O}_{V}$-module generated by the columns of $[M_{e}]$. Then $M_{e}$ is contained in a free module $F:=\mathcal{O}_{V}^{e}$. Denote by $\mathcal{R}(M)$ and $\mathcal{R}(M)$ the Rees algebras of $M$ and $M_{e}$, respectively.
@@ 3. A flatness result
Let $(X_{0},0)\subset(\mathbb{C}^{n},0)$ be a reduced complex analytic curve which is not a complete intersection. Let $(Z_{0},0)\subset(\mathbb{C}^{n},0)$ be a complete intersection which contains $(X_{0},0)$ as a subvariety. Let $(X,0)\rightarrow(Y,0)$ be a flat embedded deformation of $(X_{0},0)$ with $(Y,0)$ smooth of dimension one. Denote by $Z$ the induced deformation on $(Z_{0},0)$ from $(X,0)\rightarrow(Y,0)$. Set $W:=\overline{Z\setminus X}$ and $S:=X\times_{\mathbb{C}^{n}\times Y}W$. In this section we prove that $S$ is flat over $Y$, and in particular, $I(X_{y},W_{y}):=\sum_{s_{y}\in S_{y}}I_{s_{y}}(X_{y},W_{y})$ is constant along $Y$. In Section 4 we will show that if $(X,0)\rightarrow(Y,0)$ is a smoothing of $(X_{0},0)$, then $S$ is also reduced for a generic choice of $Z_{0}$. We begin with an algebraic result.
Lemma 3.1.
Assume $(R,\mathfrak{m})$ is a local Noetherian ring with two ideals $I$ and $J$ such that $\mathrm{depth}(R/I\cap J)\geqslant 2,\,\mathrm{depth}(R/I)\geqslant 1$ and $\mathrm{depth}(R/J)\geqslant 1$. Then $\mathfrak{m}\notin\mathrm{Ass}_{R}(R/I+J)$.
@@
Phrased in terms of local cohomology the conclusion of our statement is equivalent to $H_{\mathfrak{m}}^{0}(R/I+J)=0$. Consider the short exact sequence
[…]
Extract from the long exact sequence of local cohomology the following exact sequence
[…]
By the vanishing theorem of local cohomology (see for instance [BS12, Cor. 6.2.8]), we have
$H_{\mathfrak{m}}^{0}(R/I\oplus R/J)=0\quad\text{and}\quad H_{\mathfrak{m}}^{1}(R/I\cap J)=0.$
Therefore, $H_{\mathfrak{m}}^{0}(R/I+J)=0$. ∎
@@
In the setup above, $S$ is flat over $Y$. In particular, $I(X_{y},W_{y})$ is constant for all $y$ close to 00.

### v3

@@ Abstract.
For a reduced complex analytic curve germ we introduce an algebraic invariant we call the complete intersection discrepancy. We determine some of its basic properties. As an application we derive a multiplicity formula for the Milnor number of a reduced smoothable curve singularity generalizing a well-known formula due to Lê, Greuel and Teissier for complete intersection curves. We obtain a multiplicity characterization of Whitney equisingularity for families of locally smoothable curves.
@@
Milnor number, Euler characteristic, delta invariant, complete intersection discrepancy, conormal space, polar varieties, Hilbert–Samuel and Buchsbaum–Rim multiplicities, Gorenstein, Linkage, Whitney equisingularity
@@ 1. Introduction
Let $(X_{0},0)\subset(\mathbb{C}^{n},0)$ be a reduced complex analytic curve singularity. Let $(Z_{0},0)\subset(\mathbb{C}^{n},0)$ be a complete intersection curve which is defined by $n-1$ generic $\mathbb{C}$-linear combinations of a set of defining equations for $(X_{0},0)$. Then $(Z_{0},0)$ contains $(X_{0},0)$ as an analytic subset and $(Z_{0},0)$ is reduced (see [BLR24, Sct. 3]). Set $W_{0}:=\overline{Z_{0}\setminus X_{0}}$. Denote by $I_{0}(X_{0},W_{0})$ the intersection multiplicity of $X_{0}$ and $W_{0}$ at 00 (see [Hir57, pg. 179]). It is equal to $\dim_{\mathbb{C}}\mathcal{O}_{\mathbb{C}^{n},0}/(I_{X_{0}}+I_{W_{0}})$, where $I_{X_{0}}$ and $I_{W_{0}}$ are the ideals of $X_{0}$ and $W_{0}$ in $\mathcal{O}_{\mathbb{C}^{n},0}$, respectively. We will call $I_{0}(X_{0},W_{0})$ the complete intersection discrepancy of $X_{0}$.
We say that $(X_{0},0)$ is smoothable if there exists a flat deformation $h\colon(X,0)\rightarrow(Y,0)$ with $Y$ smooth of dimension one such that $(h^{-1}(0),0)=(X_{0},0)$ and the fiber $h^{-1}(y):=X_{y}$ is smooth for $y\neq 0$ close enough to 00.
In this paper we initiate the study of the complete intersection discrepancy. We show that it is constant under one-parameter flat deformations of $(X_{0},0)$. When $(X_{0},0)$ is smoothable we show how to compute it without having to find $I_{W_{0}}$, and using linkage we also show that the complete intersection discrepancy is an intrinsic invariant of $(X_{0},0)$ (the latter is done in full generality in [RT24]). As a first application of these results we derive in this paper a multiplicity formula for the Milnor number of a smoothable $(X_{0},0)$. In [BLR24] we use the complete intersection discrepancy to obtain a generalization of the classical genus-degree formula for plane curves. In [PR24] we use the complete intersection discrepancy to show in a new way the uniqueness of the Milnor fiber of a smoothable curve up to a diffeomorphism.
In [BG80] Buchweitz and Greuel define the Milnor number of $(X_{0},0)$ as follows. Let $n\colon\overline{(X_{0},0)}\rightarrow(X_{0},0)$ be the normalization morphism. Denote by $\omega_{X_{0},0}$ the dualizing module of Grothendieck. Denote by $d$ the composition of $\Omega_{X_{0},0}^{1}\rightarrow n_{*}\Omega_{\overline{X_{0},0}}^{1}\cong n_{*}\omega_{\overline{X_{0},0}}\rightarrow\omega_{X_{0},0}$ and the exterior derivation $\mathcal{O}_{X_{0},0}\rightarrow\Omega_{X_{0},0}^{1}$. Then the Milnor number $\mu$ of $(X_{0},0)$ is defined as
@@
In [BG80, Prp. 1.2.1] Buchweitz and Greuel showed that $\mu=2\delta-r+1$, where $\delta:=\dim_{\mathbb{C}}(\overline{\mathcal{O}_{X_{0},0}}/\mathcal{O}_{X_{0},0})$ and $r$ is the number of branches of $(X_{0},0)$. This is a formula discovered for plane curves by Milnor. When $(X_{0},0)$ is smoothable, Bassein [Bas77] proved that $\mu$ equals the first Betti number of the smoothing; in particular, for complete intersection curves $\mu$ coincides with the usual Milnor number.
@@
The main result of this paper generalizes the formula above to the case of smoothable curves by inserting in it $I_{0}(X_{0},W_{0})$ as a correction term.
@@
When $(X_{0},0)$ is a complete intersection, then $I_{W_{0}}=(1)$ and so $I_{0}(X_{0},W_{0})$ vanishes. Thus (1) recovers the Lê–Greuel–Teissier formula. In [BLR24] (1) is proved for Gorenstein curves.
@@
Whitney equisingularity. Throughout the paper when dealing with complex analytic germs we will work with suitably small representatives of these germs. Let $(X,0)\subset(\mathbb{C}^{n+k},0)$ be a reduced complex analytic space of pure dimension $k+1$, and let $(Y,0)$ be a smooth subspace of $(X,0)$ of dimension $k$ such that $X-Y$ is smooth. Choose an embedding of $(X,0)$ in $\mathbb{C}^{n+k}=\mathbb{C}^{n}\times\mathbb{C}^{k}$, so that $(Y,0)$ is represented by $0\times U$, where $U$ is an open neighborhood of 00 in $\mathbb{C}^{k}$. Let $\mathrm{pr}:\mathbb{C}^{n}\times\mathbb{C}^{k}\rightarrow\mathbb{C}^{k}$ be the projection on the second factor and let $i:\mathbb{C}^{k}\rightarrow\mathbb{C}^{n}\times\mathbb{C}^{k}$ be the inclusion $i(y)=(0,y)$. Denote by $\pi_{Y}$ be the retraction $i\circ\mathrm{pr}_{|X}$. View $X$ as the total space of the family $\pi_{Y}:(X,0)\rightarrow(Y,0)$. For each $y\in Y$ set $X_{y}:=\pi^{-1}_{Y}(y)$. We say $H$ is a tangent hyperplane at $x\in X-Y$ if $H$ is a hyperplane in $\mathbb{C}^{n+k}$ that contains the tangent space $T_{x}X$. Let $(x_{i})$ be a sequence of points from $X-Y$ and $(y_{i})$ be a sequence of points from $Y$ both converging to 00. Suppose that the sequence of secants $(\overline{x_{i}y_{i}})$ has limit $l$ and the sequence of tangent hyperplanes $\{T_{x_{i}}X\}$ has limit $T$. We say that $(X,0)\rightarrow(Y,0)$ is Whitney equisingular at 00, or that the pair of strata $(X-Y,Y)$ satisfies the Whitney conditions at 00 if $l\subset T$.
@@ 2.1. Conormal spaces
Denote $\mathrm{Image}(\delta^{*})$ by $J_{\mathrm{rel}}(X)$ and call it the relative Jacobian module of $X$. Because $\mathcal{O}_{X.0}^{p}\twoheadrightarrow I/I^{2}$, we have $(I/I^{2})^{*}\subset\mathcal{O}_{X.0}^{p}$. So $J_{\mathrm{rel}}(X)$ is contained in $\mathcal{O}_{X,0}^{p}$. Define the Rees algebra $\mathcal{R}(J_{\mathrm{rel}}(X))$ to be the subalgebra of $\mathrm{Sym}(\mathcal{O}_{X,0}^{p})$ generated by $J_{\mathrm{rel}}(X)$ placed in degree $1$. Define the relative conormal space $C_{\mathrm{rel}}(X)$ of $X$ to be the closure in $X\times\check{\mathbb{P}}^{n-1}$ of the set of pairs $(x,H)$ where $x$ is a smooth point of $X_{h(x)}$ and $H$ is a (tangent) hyperplane in $\mathbb{C}^{n}$ at $x$ containing $T_{x}X_{h(x)}$. We have
@@ 2.2. Polar varieties
Note that the proof above establishes the first equality in (3) for a smoothable $(X_{0},0)$ of arbitrary dimension. Denote by $\mu(F|X_{0})$ the Milnor number of $F|X_{0}$ (see [MvS01, Dfn. 2.1]). One can get the equality between $\mathrm{deg}_{Y}\Gamma_{\mathrm{rel}}^{1}(X)$ and $\mu+m-1$ for curves by observing that $\mathrm{deg}_{Y}\Gamma_{\mathrm{rel}}^{1}(X)=\mu(F|X_{0})$ by [MvS01, Prp. 2.2] and $\mu(F|X_{0})=\mu+m-1$ by [NT08, Dfn. 3.1 and Cor. 3.2].
@@ 2.3. Multiplicities
Let $(V,0)$ be a reduced complex analytic variety. Let $M$ be an $\mathcal{O}_{V}$-module such that its rank at the generic point of each irreducible component of $V$ is $e$. Assume $M$ is contained in a free module $\mathcal{O}_{V}^{p}$ for some $p$. Denote by $[M]\in\mathrm{Mat}(p\times c,\mathcal{O}_{V})$ a presentation matrix of $\mathcal{O}_{V}^{p}/M$. Let $A\in\text{Mat}(e\times p,\mathbb{C})$ be a generic matrix such that $[M_{e}]:=A[M]$ has rank $e$ at the generic point of each irreducible component of $(V,0)$. Let $M_{e}$ be the $\mathcal{O}_{V}$-module generated by the columns of $[M_{e}]$. Then $M_{e}$ is contained in a free module $F:=\mathcal{O}_{V}^{e}$. Denote by $\mathcal{R}(M)$ and $\mathcal{R}(M_{e})$ the Rees algebras of $M$ and $M_{e}$, respectively.
@@ 3. A flatness result
Let $(X_{0},0)\subset(\mathbb{C}^{n},0)$ be a complex analytic curve such that the origin is not an embedded point of $X_{0}$. Let $(X,0)\rightarrow(Y,0)$ be a flat deformation of $(X_{0},0)$ with $(Y,0)$ smooth of dimension one. Let $(Z,0)\rightarrow(Y,0)$ be a flat family of analytic curves such that $(Z,0)$ contains $(X,0)$ as a closed analytic subspace. Assume that $(Z,0)$ is Gorenstein (e.g. a complete intersection). Further, assume that $X_{y}$ is a union of irreducible components of $Z_{y}$ for each $y\in Y$. Set $W:=\overline{Z\setminus X}$ and $S:=X\times_{Z}W$. Define $I(X_{y},W_{y}):=\sum_{s_{y}\in S_{y}}I_{s_{y}}(X_{y},W_{y})$ where $I_{s_{y}}(X_{y},W_{y})$ is the intersection multiplicity of $X_{y}$ and $W_{y}$ at $s_{y}$.
Proposition 3.1.
In the setup above, $S$ is flat over $Y$. In particular, $I(X_{y},W_{y})$ is constant for all $y$ close to 00. Furthermore, $W_{y}=\overline{Z_{y}\setminus X_{y}}$ for all $y$.
@@
Denote by $I_{X}$ and $I_{W}$ the ideals of $X$ and $W$ in $\mathcal{O}_{Z,0}$. Let $t$ be the uniformizing parameter of $\mathcal{O}_{Y,0}$. Identify $t$ with its image in $\mathcal{O}_{Z,0}$. We have $\mathcal{O}_{Z_{0},0}=\mathcal{O}_{Z,0}/t\mathcal{O}_{Z,0}$. Denote by $I_{X_{0}}$ and $I_{W_{0}}$ the images of $I_{X}$ and $I_{W}$ in $\mathcal{O}_{Z_{0},0}$, respectively. Set $W_{0}^{\prime}:=\overline{Z_{0}\setminus X_{0}}$. Note that the primary decompositions of $I_{X_{0}}$ and $I_{W_{0}^{\prime}}$ provide a primary decomposition of $(0)$ in $\mathcal{O}_{Z_{0},0}$ via the identity $(0)=I_{X_{0}}\cap I_{W_{0}^{\prime}}$. Because $(Z,0)\rightarrow(Y,0)$ is flat and $(Z,0)$ is Cohen-Macaulay, we have $\mathrm{depth}(\mathcal{O}_{Z_{0},0})=1$. Hence the associated primes of $\mathcal{O}_{Z_{0},0}$ are all minimal. Thus, by prime avoidance, there exists $\overline{a}\in I_{X_{0}}$ such that $\overline{a}$ avoids the minimal primes of $I_{W_{0}^{\prime}}$. So if $\bar{a}\bar{b}=0$ in $\mathcal{O}_{Z_{0},0}$, then $\bar{b}\in I_{W_{0}^{\prime}}$. But $I_{X}I_{W}=0$, so $I_{X_{0}}I_{W_{0}}=\bar{0}$ in $\mathcal{O}_{Z_{0},0}$. Therefore, $I_{W_{0}}\subset I_{W_{0}^{\prime}}$.
If $S$ is empty, there is nothing to prove, so suppose otherwise. Observe that the ideal of $(S,0)$ in $\mathcal{O}_{Z,0}$ is $I_{X}+I_{W}$. We will show that $t$ is a nonzerodivisor in $\mathcal{O}_{S,0}=\mathcal{O}_{Z,0}/I_{X}+I_{W}$. Suppose there exists $z\in\mathcal{O}_{Z,0}$ such that $tz=x+w$ with $x\in I_{X}$ and $w\in I_{W}$. Denote by $\bar{x}$ and $\bar{w}$ the images of $x$ and $w$ in $\mathcal{O}_{Z_{0},0}$. We have $\bar{x}+\bar{w}=\bar{0}$. Note that $\bar{w}\in I_{W_{0}}\subset I_{W_{0}^{\prime}}$ and $\bar{x}+\bar{w}\in I_{W_{0}^{\prime}}$. So $\bar{x}\in I_{W_{0}^{\prime}}$. But $\bar{x}\in I_{X_{0}}$, so $\bar{x}=\bar{0}$. Thus we can write $x=tx_{1}$ for some $x_{1}\in\mathcal{O}_{Z,0}$. Because $X$ is flat over $Y$, we have that $t$ is a nonzerodivisor in $\mathcal{O}_{X,0}$. Hence $x_{1}\in I_{X}$. Now $tz=x+w$ gives $t(z-x_{1})=w$. Because $(Z,0)\rightarrow(Y,0)$ is flat, $t$ avoids the associated primes of $\mathcal{O}_{Z,0}$. Because $I_{W}$ is obtained from the primary decomposition of $(0)$ by removing $I_{X}$, we get that $t$ avoids the associated primes of $I_{W}$. Thus, there exists $w_{1}\in I_{W}$ such that $z-x_{1}=w_{1}$, i.e. $z=x_{1}+w_{1}\in I_{X}+I_{W}$. So, $t$ is a nonzerodivisor in $\mathcal{O}_{S,0}$.
In particular, $(S,0)$ has no vertical components and $(S,0)$ cannot be equal set-theoretically to 00. Thus $(S,0)\rightarrow(Y,0)$ is quasi-finite with nonempty fibers. Therefore, by taking sufficiently small representatives of the germs $(X,0)$ and $(Y,0)$ we can ensure that $(S,0)\rightarrow(Y,0)$ is finite (see [GrR84, Prp. 1), p. 63]). So $\mathcal{O}_{S,0}$ is a finite $\mathcal{O}_{Y,0}$-module which is torsion-free because $t$ is nonzerodivisor in $\mathcal{O}_{S,0}$. Therefore, $(S,0)\rightarrow(Y,0)$ is flat, and thus $y\rightarrow I(X_{y},W_{y})$ is constant for all $y$ close to 00.
Next we prove that the formation of $W$ specializes with passage to the fibers of $(X,0)\rightarrow(Y,0)$. We will prove it below for $y=0$, i.e.  we will show that $W_{0}=W_{0}^{\prime}$. For generic $y\neq 0$ the proof follows from the fact that the primary decomposition of the ideal $(0)$ in $\mathcal{O}_{Z,0}$ specializes with passage to generic fibers (see [BouAC, Ch. IV, §2, Prp. 11, p. 333]). We have already established above the inclusion $I_{W_{0}}\subset I_{W_{0}^{\prime}}$. Thus to prove that $W_{0}=W_{0}^{\prime}$ it’s enough to show $I_{W_{0}^{\prime}}\subset I_{W_{0}}$.
Because $\mathrm{depth}(\mathcal{O}_{X_{0},0})=1$ and $(X,0)\rightarrow(Y,0)$ is flat, we have that $\mathcal{O}_{X,0}$ is Cohen-Macaulay. By [E95, Thm. 21.23 (b) (Linkage)] $\mathcal{O}_{W,0}$ is Cohen–Macaulay because $\mathcal{O}_{Z,0}$ is Gorenstein. So $\mathrm{depth}(\mathcal{O}_{W,0})=2$. Because $\mathcal{O}_{W,0}$ is a flat $\mathcal{O}_{Y,0}$-module, we have that $t$ is a nonzerodivisor in $\mathcal{O}_{W,0}$ and $\mathrm{depth}(\mathcal{O}_{W_{0},0})=1$. In particular, the maximal ideal in $\mathcal{O}_{Z_{0},0}$ is not an associated prime of $I_{W_{0}}$. Observe that set-theoretically $Z_{0}=X_{0}\cup W_{0}$. But $S$ has no irreducible components in $Z_{0}$. This shows that $W_{0}$ and $W_{0}^{\prime}$ are set-theoretically equal. By prime avoidance we can select $r\in I_{X}$ such that $r$ avoids the minimal primes of $I_{W}$ and such that its image $\bar{r}$ in $\mathcal{O}_{Z_{0},0}$ avoids the associated primes of $I_{W_{0}}$. Thus $\bar{r}$ is a nonzerodivisor in $\mathcal{O}_{W_{0},0}$. Therefore, $(t,r)$ is a regular sequence in $\mathcal{O}_{W,0}$ and so is $(r,t)$ because $\mathcal{O}_{W,0}$ is local.
We claim that $t$ is a nonzerodivisor in $\mathcal{O}_{Z,0}/r\mathcal{O}_{Z,0}$. Suppose there exists $z\in\mathcal{O}_{Z,0}$ such that $tz\in r\mathcal{O}_{Z,0}$. Because $r\in I_{X}$ and $t$ is a nonzerodivisor in $\mathcal{O}_{X,0}$ we have $z\in I_{X}$. Also, we have $tz=0$ in $\mathcal{O}_{Z,0}/(r,I_{W})=\mathcal{O}_{W,0}/r\mathcal{O}_{W,0}$. Because $(r,t)$ is a regular sequence in $\mathcal{O}_{W,0}$ we get that $z=rz_{1}+w$ where $z_{1}\in\mathcal{O}_{Z,0}$ and $w\in I_{W}$. But $z$ and $r$ belong to $I_{X}$. Thus $w\in I_{X}$. But $I_{X}\cap I_{W}=(0)$. Therefore, $w=0$ and so $z=rz_{1}$ which establishes our claim.
Let $v\in\mathcal{O}_{Z,0}$ such that its image $\bar{v}$ in $\mathcal{O}_{Z_{0},0}$ is in $I_{W_{0}^{\prime}}$. By our choice of $r$ we have $\bar{r}\bar{v}=0$ in $\mathcal{O}_{Z_{0},0}$. Thus $rv=tz_{2}$ for some $z_{2}\in\mathcal{O}_{Z,0}$. By what we have just proved above, there exists $z_{3}\in\mathcal{O}_{Z,0}$ such that $z_{2}=rz_{3}$. Hence $r(v-tz_{3})=0$. By our choice of $r$ we must have $v-tz_{3}\in I_{W}$. Thus $\bar{v}\in I_{W_{0}}$ and so $I_{W_{0}^{\prime}}\subset I_{W_{0}}$, which finishes the proof. ∎
In our applications $(X_{0},0)$ is a reduced curve defined in $(\mathbb{C}^{n},0)$ by the vanishing of analytic functions $\tilde{f_{1}},\ldots,\tilde{f_{p}}$. Let $A\in\mathrm{Mat}((n-1)\times p,\mathbb{C})$ and let $(Z_{0},0)\subset(\mathbb{C}^{n},0)$ be the analytic set defined by the vanishing of $A(\tilde{f_{1}},\ldots,\tilde{f_{p}})^{T}$. Obviously, $(Z_{0},0)$ contains $(X_{0},0)$ set-theoretically. For a generic $A$ we show below in Prp. 3.2 that is $(Z_{0},0)$ is a complete intersection which is reduced along $(X_{0},0)$.
Fix such an $A$. Observe that any $(X,0)\rightarrow(Y,0)$ flat embedded deformation of $(X_{0},0)$ with $(Y,0)$ smooth and $(X,0)\subset(\mathbb{C}^{n}\times Y,0)$ induces an embedded deformation on the complete intersection $(Z_{0},0)$ which we denote by $(Z,0)\rightarrow(Y,0)$. Alternatively, if $(X,0)$ is cut out from $(\mathbb{C}^{n}\times Y,0)$ by $f_{1},\ldots,f_{p}$, then $(Z,0)$ is cut out in $(\mathbb{C}^{n}\times Y,0)$ by $A(f_{1},\ldots,f_{p})^{T}$. One can readily check that $(Z,0)$ is reduced along $(X,0)$ because $(Z_{0},0)$ is reduced along $(X_{0},0)$ using the Jacobian criterion for smoothness, for example.
@@
Suppose $(X_{0},0)$ is reduced and suppose $A$ is generic. Then $(Z_{0},0)$ is a complete intersection which is reduced along $(X_{0},0)$.